# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v2

@@ VII-B Proof of 1
To prove, it suffices to show that ${\mathbb{P}[\mathbf{X}_{k+1}=\bm{x}^{\mathsf{N}}(k+1)|\mathbf{X}_{k}=\bm{x}^{\mathsf{N}}(k)]=\text{Pr}_{\bm{x}^{\mathsf{N}}(k),\bm{x}^{\mathsf{N}}(k+1)}}$. That is to show that the probability the population dynamics (3) reach the state $\bm{y}^{\mathsf{N}}$ at index $k+1$ from the state $\bm{x}^{\mathsf{N}}$ at index $k$ equals the transition probability $\text{Pr}_{\bm{x}^{\mathsf{N}},\bm{y}^{\mathsf{N}}}$ in the Markov chain. This is straightforward, because at any state $\bm{x}^{\mathsf{N}}$ either the active agent switches to $\mathtt{A}$ or to $\mathtt{B}$ or sticks to her current strategy and no other case is possible. This summarizes the possible four cases in (4). In the first case, for example, the probability of an agent which belongs to the subpopulation $p$ switches her strategy from $\mathtt{B}$ to $\mathtt{A}$ equals the probability of drawing such agent ${\rho}_{p}-x^{\mathsf{N}}_{p}$ if strategy $\mathtt{A}$ is the preferred strategy of subpopulation $p$’s agents, i.e., $\smash{s^{*}(p,\bm{x}^{\mathsf{N}})=1}$, and equals zero otherwise, i.e., $\smash{s^{*}(p,\bm{x}^{\mathsf{N}})=2}$. ∎
@@ VII-C Proof of Lemma 1
The proof of Lemma 1 is inspired by the steps taken in [37, Example 4.1].
@@
By considering the sequence $\langle\frac{1}{\mathsf{N}}\rangle_{\mathsf{N}\in\mathcal{N}}$ as a vanishing sequence, it suffices to show that there exist $\langle\bm{\mathcal{V}}^{\frac{1}{\mathsf{N}}}\rangle_{{\frac{1}{\mathsf{N}}}}$ and $\langle\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ such that the conditions in Definition 4 are met. The first condition of Definition 4 can be easily verified. Let $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})$ denote the expected increment per time unit $t$ of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ when the Markov state is at $\bm{x}^{\mathsf{N}}$. Since $k=\mathsf{N}t$, there are $\mathsf{N}$ transitions per unit time resulting in $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})=\smash{\mathsf{N}\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}]}$. The $p$th element of $\bm{\nu}^{\frac{1}{\mathsf{N}}}$, $p\in[\mathsf{p}+\mathsf{p}^{\prime}]$, which is the expected increment in the subpopulation $p$, equals the summation of the multiplication of each possible change in the subpopulation $p$ with its probability. In view of (4), there are two possible changes, i.e., $\frac{1}{\mathsf{N}}$ step increment or $\frac{1}{\mathsf{N}}$ step reduction, which results in ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\mathsf{N}\big{(}\frac{1}{\mathsf{N}}(\rho_{p}-x^{\mathsf{N}}_{p})(2-s^{*}(p,\bm{x}^{\mathsf{N}}))-\frac{1}{\mathsf{N}}(x^{\mathsf{N}}_{p}(s^{*}(p,\bm{x}^{\mathsf{N}})-1))\big{)}$ and consequently, ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\rho_{p}(2-s^{*}(p,\bm{x}^{\mathsf{N}}))-x^{\mathsf{N}}_{p}$. The second condition of Definition 4 is satisfied by $\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k+1}=\mathsf{N}(\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}-\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}])$ resulting in $\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}-{\frac{1}{\mathsf{N}}}\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k+1}={\frac{1}{\mathsf{N}}}\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$. Regarding the third condition, at each $\bm{x}$, ${\bm{\nu}}^{\frac{1}{\mathsf{N}}}(\bm{x})$ should belong to the $\delta$-distance of the image of (5) over a $\delta-$neighborhood of $\bm{x}$. We know that the distance from a point to a set equals to the infimum distance of the point from the members of that set. Hence, if at each $\bm{x}$, ${\bm{\nu}}^{\frac{1}{\mathsf{N}}}(\bm{x})$ lies within the $\delta$-distance of the image of a selection from (5) over a $\delta$-neighborhood of $\bm{x}$, so does it lie for the whole differential inclusion (5). As a result, we can focus on a selection from (5) rather than the whole differential inclusion (5). We, hence, consider a selection from (5) denoted by $\dot{\bm{x}}=\bm{\nu}(\bm{x})$ where $\bm{\nu}$ is a $\mathsf{p}+\mathsf{p}^{\prime}$-dimensional function and its $p$th element ${\nu}_{p}$ equals $\rho_{p}-x_{p}$ if $\mathtt{A}$ is the preferred strategy of subpopulation $p$ and equals $-x_{p}$if the preferred strategy of subpopulation $p$ is $\mathtt{B}$. This results in ${\nu}_{p}(\bm{x})=\rho_{p}(2-s^{*}(p,\bm{x}))-x_{p}$. The expected increment per unit of time, $\bm{\nu}^{\frac{1}{\mathsf{N}}}$, is the same as the function $\bm{\nu}(\bm{x})$. As a result, the third condition is satisfied by selecting $\bm{y}$ equal to $\bm{x}$. As for the last condition, in view of Example 4.1 in [37], it can be shown that $\langle\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ is a martingale difference sequence. On the other hand, for all $k$, $|\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}|\leq\sqrt{\Sigma_{l=1}^{\mathsf{p}+\mathsf{p}^{\prime}}(1+\rho_{l})^{2}}$ and hence $\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}$ is uniformly bounded. Therefore, the conditions asserted in [37, Proposition 2.3] are satisfied and the satisfaction of the last condition is guaranteed. ∎
VII-D Proof of 2
@@
It holds that $x=\Sigma_{k=1}^{\mathsf{p}+\mathsf{p}^{\prime}}x_{k}$ and $0\leq x\leq 1$. Considering 1, the thresholds of the subpopulations are distinct, building $\mathsf{p}+\mathsf{p}^{\prime}+1$ disjoint open intervals in the unit interval, where their limit points are either 00, $1$, or the thresholds. Denote by $\mathcal{T}$ the union of the set $\{0,1\}$ with the set of all subpopulations’ thresholds. When $x\notin\mathcal{T}$, there exist $i\in\{0,1,\ldots,\mathsf{p}+1\}$ and $j\in\{0,1,\ldots,\mathsf{p}^{\prime}+1\}$ where $x\in(\tau^{\prime}_{j},\tau_{i})$ such that $\tau_{i+1}<x<\tau^{\prime}_{j+1}$. In view of (5), in this case, $\bm{\mathcal{V}}(\bm{x})$ is a singleton and for all $0<l\leq i$ anticoordinating subpopulations, $\dot{x}_{l}=\rho_{l}-x_{l}$ and for the remaining anticoordinating subpopulations, i.e., $i<l\leq\mathsf{p}$, $\dot{x}_{l}=-x_{l}$. Similarly, for coordinating subpopulations, $\dot{x}_{l}=-x_{l}$ for all $\mathsf{p}<l\leq\mathsf{p}+\mathsf{p}^{\prime}-j+1$ and $\dot{x}_{l}=\rho_{l}-x_{l}$ for all $\mathsf{p}+\mathsf{p}^{\prime}-j+1<l\leq\mathsf{p}+\mathsf{p}^{\prime}$. By summing over the elements of $\dot{\bm{x}}$, we have $\dot{x}=\Sigma_{k=1}^{\mathsf{p}+\mathsf{p}^{\prime}}\dot{x}_{k}={\eta}_{i}+{\eta}^{\prime}_{j}-x$. Therefore, for $x\in(\tau^{\prime}_{j},\tau_{i})$, $\mathcal{X}(x)=\{{\eta}_{i}+{\eta}^{\prime}_{j}-x\}$ which is the same as the last case of $\mathcal{X}$ in (9). When $x\in\mathcal{T}$, three cases may happen: $x=\tau^{\prime}_{j}$, $x=\tau_{i}$, or $x\in\{0,1\}$. When $x=\tau_{i}$, based on (5), except for $\mathcal{V}_{i}$, the elements of $\bm{\mathcal{V}}(\bm{x})$ are single-valued and contain either $\rho_{l}-x_{l}$ or $-x_{l}$. More specifically, for anticoordinating subpopulations, we have $\dot{x}_{l}=\rho_{l}-x_{l}$ for $l<i$ and $\dot{x}_{l}=-x_{l}$ for $i<l\leq\mathsf{p}$. As for coordinating subpopulations, assume that the set $\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$ is not empty and let $j=\max\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$. Then, for $l$ satisfying $\mathsf{p}+\mathsf{p}^{\prime}-j+1\leq l\leq\mathsf{p}+\mathsf{p}^{\prime}$, $\dot{x}_{l}$ will be equal to $\rho_{l}-x_{l}$, and for the remaining subpopulations will be equal to $-x_{l}$. If $\{k\in[\mathsf{p}^{\prime}]|\tau^{\prime}_{k}<\tau_{i}\}$ is empty, then $\dot{x}_{l}$ for $l>\mathsf{p}$ will be equal to $-x_{l}$. Summing over the elements of $\dot{\bm{x}}$ except the $i$th element, yields ${\eta}_{i-1}+{\eta}^{\prime}_{j}-(x-x_{i})$. As for $l=i$, we have $\dot{x}_{i}\in[-x_{i},\rho_{i}-x_{i}]$ yielding $\dot{x}\in({\eta}_{i-1}+{\eta}^{\prime}_{j}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x)$. Therefore, $\mathcal{X}(x)=[{\eta}_{i-1}+{\eta}^{\prime}_{j}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x]$, which is the same as the first case of $\mathcal{X}$ in (9). If $x=\tau^{\prime}_{j}$, a similar reasoning results in $\mathcal{X}(x)=[{\eta}_{i}+{\eta}^{\prime}_{j-1}-x,{\eta}_{i}+{\eta}^{\prime}_{j}-x]$, which is the same as the second case of $\mathcal{X}$ in (9). Finally, when $x\in\{0,1\}$, a similar reasoning yields $\dot{x}={\eta}_{\mathsf{p}}-x$ if $x=0$ and $\dot{x}={\eta}^{\prime}_{\mathsf{p}^{\prime}}-x$ if $x=1$. These two are equivalent to the first two cases of $\mathcal{X}$ in (9) as $\tau_{\mathsf{p}+1}=\tau_{0}^{\prime}=0$ and $\tau^{\prime}_{\mathsf{p}^{\prime}+1}=\tau_{0}=1$. ∎
VII-E Proof of Lemma 2

### v3

@@ VII-B Proof of 1
To prove, it suffices to show that ${\mathbb{P}[\mathbf{X}_{k+1}=\bm{x}^{\mathsf{N}}(k+1)|\mathbf{X}_{k}=\bm{x}^{\mathsf{N}}(k)]=\text{Pr}_{\bm{x}^{\mathsf{N}}(k),\bm{x}^{\mathsf{N}}(k+1)}}$. That is to show that the probability the population dynamics (3) reach the state $\bm{y}^{\mathsf{N}}$ at index $k+1$ from the state $\bm{x}^{\mathsf{N}}$ at index $k$ equals the transition probability $\text{Pr}_{\bm{x}^{\mathsf{N}},\bm{y}^{\mathsf{N}}}$ in the Markov chain. This is straightforward, because at any state $\bm{x}^{\mathsf{N}}$ either the active agent switches to $\mathtt{A}$ or to $\mathtt{B}$ or sticks to her current strategy and no other case is possible. This summarizes the possible four cases in (6). In the first case, for example, the probability of an agent that belongs to the subpopulation $p$ switches her strategy from $\mathtt{B}$ to $\mathtt{A}$ equals the probability of drawing such an agent ${\rho}_{p}-x^{\mathsf{N}}_{p}$ if strategy $\mathtt{A}$ is the preferred strategy of subpopulation $p$’s agents with current strategy $\mathtt{B}$, i.e., $\smash{s^{*}(p,\bm{x}^{\mathsf{N}},{2})=1}$, and equals zero otherwise, i.e., $\smash{s^{*}(p,\bm{x}^{\mathsf{N}},{2})=2}$. ∎
@@ VII-C Proof of Lemma 1
First, we introduce the following lemma, which facilitates the proof of Lemma 1.
Lemma A1.
Under 1, let $d$ denote the minimum distance between thresholds, i.e., $d=\min_{i,j\in[\mathsf{p}+\mathsf{p}^{\prime}]}|\tau_{i}-\tau_{j}|.$ Then, for any population size $\mathsf{N}$ satisfying $\mathsf{N}d>1$ and any $\bm{x}\in\bm{\mathcal{X}}_{ss}$, the sum $\bm{1}^{\top}\bm{x}$ belongs to at most one interval of the form $[\tau_{i},\tau_{i}+\frac{1}{\mathsf{N}})$ with $i\in[\mathsf{p}+\mathsf{p}^{\prime}]$.
@@
We assume, on the contrary, that there is also a threshold $\tau_{j}$ where $\bm{1}^{\top}\bm{x}\in[\tau_{j},\tau_{j}+\frac{1}{\mathsf{N}})$ and that $\tau_{i}\neq\tau_{j}$. Without loss of generality, let $\tau_{j}>\tau_{i}$ which results in $\tau_{j}<x<\tau_{i}+1/\mathsf{N}$ and, in turn, $\tau_{j}<\tau_{i}+1/\mathsf{N}$. However, according to the assumption $1/\mathsf{N}<d$, and we reach a contradiction. ∎
The proof of Lemma 1 is inspired by the steps taken in [38, Example 4.1].
@@
By considering the sequence $\langle\frac{1}{\mathsf{N}}\rangle_{\mathsf{N}\in\mathcal{N}}$ as a vanishing sequence, it suffices to show that there exist $\langle\bm{\mathcal{V}}^{\frac{1}{\mathsf{N}}}\rangle_{{\frac{1}{\mathsf{N}}}}$ and $\langle\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ such that the conditions in Definition 4 are met. The first condition of Definition 4 can be easily verified. Let $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})$ denote the expected increment per time unit $t$ of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ when the Markov state is at $\bm{x}^{\mathsf{N}}$. Since $k=\mathsf{N}t$, there are $\mathsf{N}$ transitions per unit time resulting in $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})=\smash{\mathsf{N}\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}]}$. The $p^{\text{th}}$ element of $\bm{\nu}^{\frac{1}{\mathsf{N}}}$, $p\in[\mathsf{p}+\mathsf{p}^{\prime}]$, which is the expected increment in the subpopulation $p$, equals the sum of the multiplication of each possible change in the subpopulation $p$ with its probability. In view of (6), there are two possible changes, i.e., $\frac{1}{\mathsf{N}}$ step increment or $\frac{1}{\mathsf{N}}$ step reduction, which results in ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\mathsf{N}\big{(}\frac{1}{\mathsf{N}}(\rho_{p}-x^{\mathsf{N}}_{p})(2-s^{*}(p,\bm{x}^{\mathsf{N}},{2}))-\frac{1}{\mathsf{N}}(x^{\mathsf{N}}_{p}(s^{*}(p,\bm{x}^{\mathsf{N}},{1})-1))\big{)}$ and consequently, ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\rho_{p}(2-s^{*}(p,\bm{x}^{\mathsf{N}},{2}))-x^{\mathsf{N}}_{p}{+x^{\mathsf{N}}_{p}\left(s^{*}(p,\bm{x}^{\mathsf{N}},2)-s^{*}(p,\bm{x}^{\mathsf{N}},1)\right)}$. The second condition of Definition 4 is satisfied by $\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k+1}=$ $\smash{\mathsf{N}(\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}-\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}])}$ resulting in $\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}-{\frac{1}{\mathsf{N}}}\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k+1}={\frac{1}{\mathsf{N}}}\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$. Now, we investigate Condition 3. To show that Condition 3 is satisfied, for a given $\delta$, we need to provide an $\epsilon_{0}$ satisfying eq. 5. We propose an $\epsilon_{0}$ which satisfies $\epsilon_{0}<\min\{d,\delta\}.$ Then, given A1, for any $\bm{x}\in\bm{\mathcal{X}}_{ss}$ there exists at most one threshold, say $\tau_{l}$, $l\in[\mathsf{p}+\mathsf{p}^{\prime}]$, that satisfies $\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\frac{1}{\mathsf{N}})$, with $\frac{1}{\mathsf{N}}\leq\epsilon_{0}.$ In what follows, we show that Condition 3 holds for all $\bm{x}\in\bm{\mathcal{X}}_{ss}$ that satisfy $\exists l\in[\mathsf{p}+\mathsf{p}^{\prime}](\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$; the other values of $\bm{x}$ can be handled similarly. Let $\varepsilon_{x}=\bm{1}^{\top}\bm{x}-\tau_{l}$. Then given $\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$, we have $\varepsilon_{x}<\epsilon_{0}$ and, in turn, $\varepsilon_{x}<\delta.$ Consider the point $\bm{y}$ whose all elements, except some element $j$, are equal to the corresponding elements of $\bm{x}$, and the $j^{\text{th}}$ element, $y_{j}$, equals $x_{j}-\varepsilon_{x}$. The term $|\bm{x}-\bm{y}|$ then equals $\varepsilon_{x}$, and hence the point $\bm{y}$, which satisfies $\bm{1}^{\top}\bm{y}=\tau_{l}$, belongs to the $\delta$-neighborhood of $\bm{x}$, i.e., $|\bm{x}-\bm{y}|<\delta$. For all $p\neq l$, we then have $\mathcal{V}_{p}({\bm{y}})=2(\rho_{p}-s^{*}(p,{\bm{y}},2))-{y}_{p}$ and for $p=l$, $\mathcal{V}_{p}=[-{y}_{p},\rho_{p}-{y}_{p}]$. On the other hand, the relation $\exists l\in[\mathsf{p}+\mathsf{p}^{\prime}](\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$ implies that ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}})$ for $p=l$ equals $\rho_{l}-2x^{\frac{1}{\mathsf{N}}}_{l}$ (resp. 00) if $l$ corresponds to a coordinating (resp. an anticoordinating) subpopulation. For $p\neq l$, ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}})$ equals $2(\rho_{p}-s^{*}(p,\bm{x},2))-x_{p}$, and it can be easily shown that it is also equal to the unique member of $\mathcal{V}_{p}(\bm{x})$ defined in (7). By $\bm{z}=\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$ and choosing $\bm{v}$ to be a member of $\bm{\mathcal{V}}({\bm{y}})$ with $v_{l}({\bm{y}})=\rho_{l}-2{y}_{l}$ (resp. $v_{l}({\bm{y}})=0$) if $l$ corresponds to a coordinating (resp. an anticoordinating) subpopulation, the term $|\bm{z}-\bm{v}|$ equals $|\bm{v}({\bm{y}})-\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})|=|{\bm{y}}-\bm{x}|$ which is less than $\delta$. This implies the satisfaction $\inf_{\bm{v}\in\bm{\mathcal{V}}(\bm{y})}\!|\bm{z}-\bm{v}|<\delta$ and, in turn, Condition 3. As for the last condition, in view of Example 4.1 in [38], it can be shown that $\langle\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ is a martingale difference sequence. On the other hand, for all $k$, $|\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}|\leq\sqrt{\Sigma_{l=1}^{\mathsf{p}+\mathsf{p}^{\prime}}(1+\rho_{l})^{2}}$ and hence $\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}$ is uniformly bounded. Therefore, the conditions asserted in [38, Proposition 2.3] are satisfied and the satisfaction of the last condition is guaranteed. ∎
VII-D Proof of Lemma 1 when agents include themselves
As noted in Remark 3, in some set-ups, agents take into account their current strategies when they want to determine their proffered strategies. In this case, the update rule (1) changes to
$\mathtt{s}_{i}(k+1)=\begin{cases}\mathtt{A}&\text{if }x^{\mathsf{N}}(k)\geq{\tau(i)},\\ \mathtt{B}&\text{otherwise},\end{cases}$
and the update rule (2) changes to
$\mathtt{s}_{i}(k+1)=\begin{cases}\mathtt{A}&\text{if }x^{\mathsf{N}}(k)\leq{\tau(i)},\\ \mathtt{B}&\text{otherwise},\end{cases}$
where $x^{\mathsf{N}}$ is the population proportion of $\mathtt{A}$-players. Accordingly, the function $s^{*}(\cdot,\cdot,\cdot)$, defined in Definition 1 and used in Equation 6, changes to
${\hat{s}}(p,\bm{x}^{\mathsf{N}})=$
$\begin{cases}{1}&\text{if }(x^{\mathsf{N}}\leq{\tau_{p}}\text{ and }p\leq\mathsf{p})\text{ or }(x^{\mathsf{N}}\geq\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}\text{ and }p>\mathsf{p}),\\ {2}&\text{if }(x^{\mathsf{N}}>{\tau_{p}}\ \text{ and }p\leq\mathsf{p})\text{ or }(x^{\mathsf{N}}<\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}\text{ and }p>\mathsf{p}).\end{cases}$
If Lemma 1 remains true, then the main results of the papers are also valid. Condition 1 is already satisfied. Let $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})$ denote the expected increment per time unit $t$ of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ when the Markov state is at $\bm{x}^{\mathsf{N}}$, i.e., $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})=\smash{\mathsf{N}\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}]}$. The $p$th element of $\bm{\nu}^{\frac{1}{\mathsf{N}}}$, $p\in[\mathsf{p}+\mathsf{p}^{\prime}]$, reads as ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\rho_{p}(2-\hat{s}(p,\bm{x}^{\mathsf{N}}))-x^{\mathsf{N}}_{p}$. Accordingly, Conditions 2 and 4 can be easily verified. Regarding Condition 3, For each given $\delta$, we need to find an $\epsilon_{0}$ which satisfies (5). In what follows, we show that (5) is satisfied for any arbitrarily chosen $\epsilon_{0}$. We consider a selection from (7) denoted by $\dot{\bm{x}}=\bm{\nu}(\bm{x})$ where $\bm{\nu}$ is a $(\mathsf{p}+\mathsf{p}^{\prime})$-dimensional function and its $p$th element ${\nu}_{p}$ equals $\rho_{p}-x_{p}$ if $\mathtt{A}$ is the preferred strategy of subpopulation $p$ and equals $-x_{p}$ if the preferred strategy of subpopulation $p$ is $\mathtt{B}$. This results in ${\nu}_{p}(\bm{x})=\rho_{p}(2-\hat{s}(p,\bm{x}))-x_{p}$. The expected increment per unit of time, $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$, is the same as the function $\bm{\nu}(\bm{x})$. As a result, equation (5) and, in turn, Condition 3 are satisfied by selecting $\bm{y}$ equal to $\bm{x}$, regardless of the value of $\epsilon_{0}$. This implies that Lemma 1 holds.
VII-E Proof of 2